Ultrasonic waves of frequency 4.5 MHz are used to detect tumour in soft tissues. The speed of sound in tissue is 1.5 km/s and that in air is 340 m/s. Find the wavelength of this ultrasonic wave in air and in tissue.

Grade:12

1 Answers

Arun
6 years ago
Given V base air = 340 m/s , n = 4.5 ×10^6 Hz ⇒λ base air = (340 / 4.5) × 10^–6 = 7.36 × 10¬^–5 m. b) V base tissue = 1500 m/s ⇒ λ base t = (1500 / 4.5) × 10^–6 = 3.3 × 10^–4 m